OpenStudy (anonymous):

When the quadratic equation y = (x - 3)(2x + 7) is written in standard form, what is the value of the coefficient "b"?

OpenStudy (anonymous):

Do you know how to use the "FOIL" rule. Standard qudratic notation is \[a x^{2}+bx+c\]

OpenStudy (inkyvoyd):

HINT: standard form

OpenStudy (inkyvoyd):

ax^2+bx+c=y

OpenStudy (inkyvoyd):

you are looking for b, as stated.

OpenStudy (anonymous):

yes, expand the brackets, and look for the coefficient of x SPOILER: (x -3)(2x+7) = x(2x+7) - 3(2x + 7) = 2x^2 + 7x - 6x - 21 = 2x^2 + x - 21